# VidEdit: Zero-shot and Spatially Aware Text-driven Video Editing

### v1

@@ 4.2 State-of-the-art comparison
To further analyze the fine-grained editing capacity of our method while preserving the original video content, we display for all baselines in Fig. 4, their local $\mathcal{A}_{\text{\scriptsize{Frame}}}$ score computed within a ground truth mask compared to an outer LPIPS metric (denoted O-LPIPS) computed on the invert of the mask [^2] The size of each dot is proportional to the local frame accuracy’s standard deviation.. We see that VidEdit reaches a very good local frame accuracy, even outperforming Text2Live. Morevover, VidEdit shows a huge improvement on the O-LPIPS metric compared to the baselines, including Text2Live (3 vs 8), showing clearly a better preservation of out-of-interest regions.
@@
Qualitative results. We show in Fig. 5 a visual comparison against the baselines to qualitatively assess the improvement brought out by our method. We can see that VidEdit performs fine-grained editing while perfectly preserving out-of-interest regions. In comparison to other baselines, the edits generated are more visually appealing and realistic. For example, VidEdit obtains a frame accuracy ($\mathcal{A}_{\text{\scriptsize{Frame}}}$) and prompt consistency ($\mathcal{C}_{\text{\scriptsize{Prompt}}}$) scores of 26.5 and 30 respectively compared to Text2Live which reaches 26 and 35 respectively. However, we can see that Text2Live’s scores do not automatically translate into high-quality edits it often struggles to render detailed textures precisely localized on targeted regions. For example, ice creams are poorly rendered and some untargeted areas are being altered. Regarding Tune-a-Video and Pix2Video baselines, the methods are unable to generate a faithful edit at the exact location and completely degrade the original content. Despite relatively high frame consistency scores in this video (96% for Tune-a-Video and 89% for Pix2Video vs 97.5% for VidEdit), noticeable flickering artifacts undermine the video content. On the other hand, naive frame-wise application of image-to-image translation methods also leads to temporally inconsistent results. For example, SDEdit is unable to both generate a faithful edit and to preserve the original content as it inherently faces a trade-off between the two. Other visual comparisons are shown in Supplementary C.
@@
4.3 Model Analysis
Ablations. We perform ablation studies to demonstrate the importance of our conditional controls once we map the edits back to the original image space. Tab. 2 compares the performance of our editing pipeline with both instance mask segmentation and HED edge conditioning against scenarios where these controls are disabled.
In the case where no conditional control is passed on to the model, we observe a substantial drop in semantic metrics as the model generates edits at random locations in the atlas whose shapes don’t match the structure of the target object. The introduction of edge conditioning without spatial awareness is quite similar to the previous case with the difference that the model tries to locally match the control information. This results in slightly better semantic results and similarity metrics than with no edge control. Finally, blending an edit with mask control without taking structure conditioning into account generates an edit at the right location but that is semantically incoherent once mapped back to the original images. Yet, this scenario achieves a decent prompt consistency as the objects still correspond to the target text query. We provide a visual illustration of this ablation study in Supplementary B.
Figure 6: VidEdit behavior wrt. different $\lambda$ and $\rho$ values.
On the left panel, we can see that for $\lambda$ values lower than 0.4, the edge conditioning is not strong enough to guide the edits toward a plausible output on the video frames. This phenomenon is illustrated in Supplementary B. On the contrary, for strength values larger than 1.2, the conditioning weighs too much on the model and hinders its ability to generate faithful edits. As expected, we notice that the local LPIPS decreases as the edge conditioning increases. While the decreasing rate is substantial between 0 and 1, the marginal gain diminishes for larger values. Overall, setting the HED strength between 0.8 and 1.2 robustly enables to both perform faithful edits and preserve the original content. On the right panel, we see that both local CLIP score and LPIPS increase with the noising ratio. Indeed, for a null $\rho$ value, the region is reconstructed from the atlas, nearly identically to the original, and is then rewarded a low LPIPS. However, as no modification has been performed, the patch does not match the target text query and gets a lower CLIP score. As the noising ratio increases, the region deviates more from the input but also better matches the target edit. Note that for a $\rho$ value of 100%, the local LPIPS is constrained below 19, which still indicates a low disparity with the original image.
Figure 7: Texture diversity. We edit each video four times with the same input text query. Compared to Text2Live, our method is able to synthesize more diverse samples in much less time.
Diversity. Finally, we illustrate in Fig. 7 the capacity of VidEdit to produce various and sundry video edits from a unique pair (video; target text query). In contrast, the randomness in Text2Live’s training process only comes from the generator’s weights initialization. As a result, method converges towards a unique solution and thus shows poor diversity in the generated samples.
5 Conclusion & Discussion
We introduced VidEdit, a lightweight algorithm for zero-shot semantic video editing based on latent diffusion models. We have shown experimentally that this approach conserves more appearance information from the input video than other diffusion-based methods, leading to lighter edits. Nevertheless, the approach has a few limitations. Common with Kasten et al. (2021), the capacity of the MLP mapping networks decreases for complex videos involving rapid movements and very long-term videos. Since our method relies on the quality of such atlas representations, one possible way to expand the scope of possible video edits would be to strengthen and robustify the neural layered atlases construction approach. Finally, video editing raises several ethical challenges that we discuss here. Open-source diffusion models are trained on large amounts of data and inherit their biases. Therefore, the use of such models can raise ethical issues, whether the target text prompt is intentionally harmful or not. In addition, video editing algorithms could be used with harmful intent such as harassment or propagating fake news. To mitigate potential misuse, the Stable Diffusion model is released under a license focused on ethical and legal use, stating explicitly that users “must not distribute harmful, offensive, dehumanizing content or otherwise harmful representations of people or their environments, cultures, religions, etc. produced with the model weights”.
Appendix A Experimental Setup
A.1 Semantic Metrics
Prompt Consistency. $\mathcal{C}_{\text{\scriptsize{Prompt}}}$ measures the average CLIP similarity between a target text query and each video frame, for all edited videos. For a unique pair (image; caption) the CLIP similarity writes: $CLIPScore(I,C)=\max(100*cos(E_{I},E_{C}),0)$ with $E_{I}$ the visual CLIP embedding for an image $I$, and $E_{C}$ the textual CLIP embedding for a caption $C$. The prompt consistency score then writes :
$\mathcal{C}_{\text{\scriptsize{Prompt}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{i=1}^{F_{k}}CLIPScore(V_{i}^{k},C^{k})$
where $V_{i}^{k}$ is the $i$th frame of video $k$, $C^{k}$ the edited caption of video $k$, $F_{k}$ the number of frames in video $k$ and $N$ the number of edited videos.
Frame Accuracy. $\mathcal{A}_{\text{\scriptsize{Frame}}}$ corresponds to average percentage of edited frames that have a higher CLIP similarity with the target text query than with their source caption, for all edited videos.
Directional Similarity. $\mathcal{S}_{\text{\scriptsize{Dir}}}$ quantifies how closely the alterations made to an original image align with the changes between a source caption and a target caption. For a single edited frame the similarity score writes: $SIMScore(I_{t},I_{s},C_{t},C_{s})=100*cos(E_{I_{t}}-E_{I_{s}},E_{C_{t}}-E_{C_{s}})$ with $E_{I_{t}}$ and $E_{I_{s}}$ the visual CLIP embeddings of the edited image $I_{t}$ and source image $I_{s}$, and $E_{C_{t}}$ and $E_{C_{s}}$ the textual CLIP embeddings of the edited caption $C_{t}$ and source caption $C_{s}$. The global directional similarity score then writes:
$\mathcal{S}_{\text{\scriptsize{Dir}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{i=1}^{F_{k}}SIMScore\left(\prescript{i}{}{V_{t}^{k}},\prescript{i}{}{V_{s}^{k}},C_{t}^{k},C_{s}^{k}\right)$
with $\prescript{i}{}{V_{t}^{k}}$ and $\prescript{i}{}{V_{s}^{k}}$ the $i$th frame of the $k$th edited video and its source counterpart respectively and $C_{t}^{k},C_{s}^{k}$ their associated captions.
Frame Consistency. $\mathcal{C}_{\text{\scriptsize{Frame}}}$ is used to assess the temporal coherence of a video. It measures the CLIP similarity between consecutive video frames.
Aggregate Score. This metric synthesizes in a single score the overall performance of each model on semantic and similarity aspects, relatively to the best baseline. When dealing with metrics where a higher value is considered preferable, a coefficient in the aggregate score is computed as: max$(\mathcal{S}_{i})$/$\mathcal{S}_{i}^{j}$ i.e. the best score for metric $i$ divided by the score of baseline $j$ for metric $i$. When the objective is to minimize the metric, we take the inverse value. The minimal and best aggregate score for each aspect is 3, as we aggregate three semantic or similarity scores.
A.2 Similarity Metrics
Regarding content preservation, we have chosen three metrics that operate on different feature spaces in order to capture a rich description of perceptual similarities. LPIPS Zhang et al. [2018] operates on the deep feature space of a VGG network and has been shown to match human perception well. HaarPSI Reisenhofer et al. [2018] operates on a wavelet space and shows the highest correlation with human opinion. Finally, PSNR assesses the distance between images directly at a pixel level. These metrics are extensively described in the literature and we refer to it for further details.
Appendix B Ablation Visualization
Fig. 8 illustrates the ablation study we led in Tab. 2. When VidEdit receives both conditional controls, it produces high quality results. Conversely, when these controls are deactivated, the model is free to perform edits at random locations in the atlas, resulting in uninterpretable visual outcomes. Enabling only the edge conditioning yields similar results, with the difference that the model attempts to locally match inner and outer edges. Finally, the sole use of a mask allows to perform edits at the correct locations, but that are semantically absurd once mapped back to the image space.
Figure 8: Ablation visualization.
Appendix C Additional Results
C.1 VidEdit samples
Figure 9: Additional VidEdit sample results.
C.2 Baselines Comparison
@@
Appendix D Blending Effect
@@
Appendix E Editing Time

### v2

@@ 3 Experiments
To further analyze the fine-grained editing capacity of our method while preserving the original content, we display a local $\mathcal{A}_{\text{\scriptsize{Frame}}}$ score computed within a ground truth mask compared to an outer LPIPS metric (denoted O-LPIPS) computed on the invert of the mask. The size of each dot is proportional to the standard deviation of the local frame accuracy. VidEdit reaches a very good local $\mathcal{A}_{\text{\scriptsize{Frame}}}$ while showing a huge improvement on the O-LPIPS metric compared to all baselines. This clearly indicates a better preservation of out-of-interest regions.
Ablations. To highlight the importance of our conditional controls, we present a performance comparison in Tab. 2, evaluating our pipeline with mask segmentation and HED conditioning enabled versus scenarios where these controls are turned off. Without any control input, we observe a substantial drop in semantic metrics as the model generates edits randomly across the atlas with shapes that do not align with the target object’s structure.
Introducing edge conditioning without spatial awareness closely resembles the previous scenario, with the distinction that the model attempts to achieve local alignment with the control information. This leads to modestly improved semantic and similarity scores. Finally, blending an edit with mask control disregarding edge conditioning produces an edit at the correct location but that is semantically incoherent once mapped back to the original images. Yet, this scenario achieves a decent prompt consistency as the objects still correspond to the target text query. Supplementary C provides a visual illustration of this ablation study.
Figure 5: VidEdit and Text2Live sample diversity.
Diversity. Finally, we illustrate in Fig. 5 the ability of VidEdit to produce diverse edits from a unique pair (video; text query). Unlike Text2Live which converges towards a unique solution, our method is able to synthesize more diverse and convincing samples in much less time.
Future Work. VidEdit is a lightweight algorithm for zero-shot semantic video editing based on latent diffusion models. Experimentally this approach preserves more appearance information from the input video than other diffusion-based methods, leading to lighter edits. Since our method relies on the quality of such atlas representations, one way to expand the scope of possible video edits would be to strengthen and robustify the neural layered atlases construction approach.
Supplementary Material
[…]
Appendix A Experimental Setup
Dataset. Following [Bar-Tal et al., 2022, Wu et al., 2022, Qi et al., 2023], we evaluate our approach on videos from DAVIS dataset [Pont-Tuset et al., 2017] resized at a $768\times 432$ resolution. The length of these videos ranges from 20 to 70 frames. To automatically create edit prompts, we use a captioning model [Li et al., 2022] to obtain descriptions of the original video content and we manually design 4 editing prompts for each video.
VidEdit setup. Our experiments are based on latent diffusion models [Rombach et al., 2022]. We use a version of stable diffusion trained with HED edge detection [Zhang and Agrawala, 2023] at a $512\times 512$ resolution on LAION-5B dataset [Schuhmann et al., 2022]. We choose Mask2former [Cheng et al., 2022] as our instance segmentation network. To edit an atlas, we sample pure Gaussian noise (i.e. $\rho=1$) and denoise it for 50 steps with DDIM sampling and classifier-free guidance [Ho and Salimans, 2022]. For a single 70 frames video, it takes $\sim 15$ seconds to edit a $512\times 512$ patch in an atlas and $\sim 1$ minute to reconstruct the video with the edit layer on a NVIDIA TITAN RTX. We set up the HED strength $\lambda$ to 1 by default.
Baselines. We compare our method with two text-to-image frame-wise editing approaches and three text-to-video editing baselines.
  1. 1.
    SDEdit [Meng et al., 2021] is a framewise zero-shot editing approach that corrupts an input frame with noise and denoise it with a target text prompt.
  2. 2.
    ControlNet [Zhang and Agrawala, 2023] performs frame-wise editing with an external condition extracted from the target frame.
  3. 3.
    Text2Live [Bar-Tal et al., 2022] is a Neural Layered Atlas (NLA) based method that trains a generator for each text query to optimize a CLIP-based loss.
  4. 4.
    Tune-a-Video (TAV) [Wu et al., 2022] fine-tunes an inflated version of a pre-trained diffusion model on a video to produce similar content.
  5. 5.
    Pix2Video [Ceylan et al., 2023] uses a structure-guided image diffusion model to perform text-guided edits on a key frame and propagate the changes to the future frames via self-attention feature injection.
Metrics. A video edit is expected to (1) be temporally consistent (Temporal), (2) faithfully render a target text query (Semantics) and (3) preserve out-of-interest regions unaltered (Similarity). To evaluate CLIP based metrics, we used CLIP ViT-L/14 [Radford et al., 2021].
  1. $\bullet$
    Temporal
    1. $\cdot$
      Frame Consistency ($\mathcal{C}_{\text{\scriptsize{Frame}}}$). Assesses the temporal coherence of a video. It measures the CLIP similarity between consecutive video frames.
      $\mathcal{C}_{\text{\scriptsize{Frame}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}-1}CLIPScore(\bar{I}_{j}^{k},\bar{I}_{j+1}^{k})$
      where $\bar{I}_{j}^{k}$ is the $j$-th frame of edited video $k$, $F_{k}$ the number of frames in video $k$ and $N$ the number of edited videos.
  2. $\bullet$
    Semantics
    1. $\cdot$
      Prompt Consistency ($\mathcal{C}_{\text{\scriptsize{Prompt}}}$). Measures the average CLIP similarity between a target text query and each video frame.
      $\mathcal{C}_{\text{\scriptsize{Prompt}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}}CLIPScore(\bar{I}_{j}^{k},\bar{C}^{k})$
      with $\bar{C}^{k}$ the edited caption of video $k$.
    2. $\cdot$
      Frame Accuracy ($\mathcal{A}_{\text{\scriptsize{Frame}}}$). Corresponds to the average percentage of edited frames that have a higher CLIP similarity with the target text query than with their source caption.
      $\mathcal{A}_{\text{\scriptsize{Frame}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}}\mathds{1}\{CLIPScore(\bar{I}_{j}^{k},\bar{C}^{k})>CLIPScore(\bar{I}_{j}^{k},C^{k})\}\times 100$
    3. $\cdot$
      Directional Similarity ($\mathcal{S}_{\text{\scriptsize{Dir}}}$). Quantifies how closely the alterations made to an original image align with the changes between the source and target caption. For a single frame the similarity score writes: $SIMScore(\bar{I}_{j}^{k},I_{j}^{k},\bar{C}^{k},C^{k})=100*cos(E_{\bar{I}_{j}^{k}}-E_{I_{j}^{k}},E_{\bar{C}_{j}^{k}}-E_{C_{j}^{k}})$.
      $\mathcal{S}_{\text{\scriptsize{Dir}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}}SIMScore\left(\bar{I}_{j}^{k},I_{j}^{k},\bar{C}^{k},C^{k}\right)$
  3. $\bullet$
    Similarity
    Regarding content preservation, we have chosen three metrics that operate on different feature spaces in order to capture a rich description of perceptual similarities.
    1. $\cdot$
      LPIPS [Zhang et al., 2018] operates on the deep feature space of a VGG network and has been shown to match human perception well.
    2. $\cdot$
      HaarPSI [Reisenhofer et al., 2018] While LPIPS evaluates the perceptual similarity between two images in a deep feature space, HaarPSI performs a Haar wavelet decomposition to assess local similarities.
    3. $\cdot$
      PSNR measures the distance with an original image in the pixel space.
    These metrics are extensively described in the literature and we refer to it for further details.
  4. $\bullet$
    Aggregate Score.
    This metric synthesizes in a single score the overall performance of each model on semantic and similarity aspects, relatively to the best baseline. When dealing with metrics where a higher value is considered preferable, a coefficient in the aggregate score is computed as: max$(\mathcal{S}_{i})$/$\mathcal{S}_{i}^{j}$ i.e. the best score for metric $i$ divided by the score of baseline $j$ for metric $i$. When the objective is to minimize the metric, we take the inverse value. The minimal and best aggregate score for each aspect is 3, as we aggregate three semantic or similarity scores.
Appendix B Qualitative Results
B.1 VidEdit samples
Figure 6: Additional VidEdit sample results.
B.2 Comparison against baselines
@@
We show in Fig. 7 a visual comparison against the baselines to qualitatively assess the improvement brought out by our method. We can see that VidEdit performs fine-grained editing while perfectly preserving out-of-interest regions. In comparison to other baselines, the edits generated are more visually appealing and realistic. For example, VidEdit obtains a frame accuracy ($\mathcal{A}_{\text{\scriptsize{Frame}}}$) and prompt consistency ($\mathcal{C}_{\text{\scriptsize{Prompt}}}$) scores of 26.5 and 30 respectively compared to Text2Live which reaches 26 and 35 respectively. However, we can see that Text2Live’s scores do not automatically translate into high-quality edits it often struggles to render detailed textures precisely localized on targeted regions. For example, ice creams are poorly rendered and some untargeted areas are being altered. Regarding Tune-a-Video and Pix2Video baselines, the methods are unable to generate a faithful edit at the exact location and completely degrade the original content. Despite relatively high frame consistency scores in this video (96% for Tune-a-Video and 89% for Pix2Video vs 97.5% for VidEdit), noticeable flickering artifacts undermine the video content. On the other hand, naive frame-wise application of image-to-image translation methods also leads to temporally inconsistent results. For example, SDEdit is unable to both generate a faithful edit and to preserve the original content as it inherently faces a trade-off between the two.
@@
Appendix C Ablation Visualization
Fig. 9 illustrates the ablation study we led in Tab. 2. When VidEdit receives both conditional controls, it produces high quality results. Conversely, when these controls are deactivated, the model is free to perform edits at random locations in the atlas, resulting in uninterpretable visual outcomes. Enabling only the edge conditioning yields similar results, with the difference that the model attempts to locally match inner and outer edges. Finally, the sole use of a mask allows to perform edits at the correct locations, but that are semantically absurd once mapped back to the image space. We provide additional qualitative results in this anonymous git repository https://anonymous.4open.science/r/videdit-videos-1BDC
Figure 9: Ablation visualization.
Appendix D Model Analysis
Figure 10: VidEdit behavior wrt. different $\lambda$ and $\rho$ values.
On the left panel, we can see that for $\lambda$ values lower than 0.4, the edge conditioning is not strong enough to guide the edits toward a plausible output on the video frames. This phenomenon is illustrated in Supplementary C. On the contrary, for strength values larger than 1.2, the conditioning weighs too much on the model and hinders its ability to generate faithful edits. As expected, we notice that the local LPIPS decreases as the edge conditioning increases. While the decreasing rate is substantial between 0 and 1, the marginal gain diminishes for larger values. Overall, setting the HED strength between 0.8 and 1.2 robustly enables to both perform faithful edits and preserve the original content. On the right panel, we see that both local CLIP score and LPIPS  increase with the noising ratio. Indeed, for a null $\rho$ value, the region is reconstructed from the atlas, nearly identically to the original, and is then rewarded a low LPIPS. However, as no modification has been performed, the patch does not match the target text query and gets a lower CLIP score. As the noising ratio increases, the region deviates more from the input but also better matches the target edit. Note that for a $\rho$ value of 100%, the local LPIPS is constrained below 19, which still indicates a low disparity with the original image.
Appendix E Blending Effect
@@
Appendix F Editing Time